OpenStudy (anonymous):

HELP:what is a correct comparison of the fractions? 7/12 and 11/20 A.7/12 < 11/20 B.7/12 > 11/20 C.7/12 =11/20

OpenStudy (jadeishere):

7/12 = 11/20? 7 * 20 = 140, 12 * 11 = 132 so does 132 = 140? No, so not C 7/12 > 11/20 is true, so B

OpenStudy (anonymous):

Yea I think B